OpenStudy (joachim):

I need help with this optimization question:

OpenStudy (joachim):

A poster is to have left and right hand margins of 6cm and the top and bottom margins are to be 4cm. The area of the inner printed part is to be 2400cm^2. Find the width of the poster satisfying these requirements and having the smallest possible total area.

OpenStudy (dumbcow):

Area = L*W inner dimensions: L-8 and W-12 --> (L-8)(W-12) = 2400 solve for L L = 2400/(W-12) + 8 substitute into Area equation, then take derivative and set equal to 0 solve for W

OpenStudy (joachim):

thanks a lot

OpenStudy (dumbcow):

no problem